A penny is dropped from the top of a 900 m tower how long does it take to hit the ground

Answer by scott8148(6628)

starting from rest, with a uniform acceleration (gravity) ___ 900 = (1/2) * g * t^2

the gravitational acceleration (g) is 9.8 m/s^2 ___ t is in seconds

900 / 4.9 = t^2

use a calculator to find t